This is the check list for experts on D2n data checker shift.
Notice : Please add items of your specialty field, so every one will do the check
Beamline, beam charge, beam polarization
- BPM A and B (x,y) positions
- Check the front room white board for default value
- Beam charge asymmetry is small (less than 1000ppm)
- Check end-of-run
- Check beam charge asymmetry stripchart
- Check that the beam charge at end-of-run makes sense ?
- All BCM charge reading should be consistent
- Example of Good Run:
APPROXIMATE BCM CHARGES (C) BCM u1 [ 0]:0.03799 BCM u3 [ 0]:0.03798 BCM u10[ 0]:0.0336 BCM u1 [++]:0.009217 BCM u3 [++]:0.009211 BCM u10[++]:0.008148 BCM u1 [--]:0.009464 BCM u3 [--]:0.009462 BCM u10[--]:0.008371 BCM u1 [+-]:0.009214 BCM u3 [+-]:0.009211 BCM u10[+-]:0.008148 BCM u1 [-+]:0.009466 BCM u3 [-+]:0.009465 BCM u10[-+]:0.008373 BCM d1 [ 0]:0.03786 BCM d3 [ 0]:0.03787 BCM d10[ 0]:0.03787 BCM d1 [++]:0.009183 BCM d3 [++]:0.009198 BCM d10[++]:0.009188 BCM d1 [--]:0.009428 BCM d3 [--]:0.009432 BCM d10[--]:0.009432 BCM d1 [+-]:0.009182 BCM d3 [+-]:0.009185 BCM d10[+-]:0.009185 BCM d1 [-+]:0.009431 BCM d3 [-+]:0.009434 BCM d10[-+]:0.009436
- Check scaler rate screen, pre-scale factors agree with earlier runs. example of normal rate @ 15uA prodction
- Check online CODA rate and CODA dead time. (Run scaler script, verify deadtime)
- Check start-of-run input. Target type, beam half wave plate, etc.
- Restart Auxiliary DAQs (HAPPEX and both Comptons) every two hours or so.
- Look at shift worker's run-sheet records.
- Confirm CODA not crashed, check DAQ-configuration:
- LHRS: Lefthrs
- BigBite: bigbite_buf
- Check Halog entries.
- Check target polarization history.
- Go To Target Status History
- Check all HV on. BB chamber HV on before run starts.
- Example Start Run with All HV On
- Check Section
- Left S2 HV Left/Right PMTs
- Left S1/Cerenkov/A1 Cerenkov HV
- Left Pion Rejector Layer 1/2 HV
- Map BB_SC_R Measured Voltage/Current
- Map BB_DC Measured Voltage/Current
- Map BB_PS_R Measured Voltage/Current
- Map BB_SC_L Measured Voltage/Current
- Map BB_PS_L Measured Voltage/Current
- Map BB_SH Measured Voltage/Current
- Halog left arm VDC efficiency plots once a shift
- Do Compton monitoring
- Look at Replays (See below)
- Update d2n Run Database
- Go to /adaq/e06014/expert_shift
- Run ./setAnalysisInfo.pl
- It will tell you what command line options are available
Run show_all(<run number>) script after shift worker finish a replay. It will pop up a set of plots for expert, which would take ~10min to check through.
- If the replay number is less than 20000, the analysis is for BigBite.
- If the replay number is greater than 20000, the analysis is for LHRS.
- Be sure to run show_all(<run number>) for each.
- Once per left HRS kinematic, do a full physics replay / replay_phys_L(run#,-1)
- When prompted change the root file name from the default, so as not to overwrite the online physics replay root file:
- Please post the "PID Detectors (Good Electrons)" page into the Halog indicating the number of good electrons from the replayed run.
A list of minimum check items : (Item / related online plots)
- Check left-HRS VDC efficiency / online("detectorL",run#)
- Check left-HRS VDC time / online("detectorL",run#)
- Check BB chamber wire map / online("detectorBB",run#)
- Check BB shower vs preshower ADC plot, threshold not changed / online("physics",run#)
- Check scaler consistency / online("scalar",run#)
- Spot++ / equavlent plots (raster current, BPMs) are included in online("detectorL",run#)
- Check syncroc signals ("Sync check 1 (straight lines)" plot from online BB plots. If there are any gaps or discontinuities in the straight line, you or the shift worker should make a HALOG post.
- Check L1A vs. clock. You can do this with the macro CheckL1A.C, which is located in the onlana directory:
.L CheckL1A.C CheckL1A(run#)
You're looking for a straight, flat line. It's okay if the first data point is a bit off (as in Fig. 2 here), or if there's a beam trip (as in the structures around 200 and 300 here; note that l1a drops all the way to zero). What's NOT ok is weird behavior that takes steady-beam running away from a straight line. HALOG plots like this.
Hint : You can try to do_all() the golden run to find out what a normal plots should look like
A list of known problems / possible bad runs
- BB chamber HV not on all the time.
- Left-arm VDC gas problem, low efficiency.
- Left-arm TDC shifts, symptom is current run and golden run shows significant difference in TDC times.
- Stop and start the three auxiliary DAQs (Compton Counting, Compton FADC and HAPPEX) roughly every two hours. These are run through a VNC connection to the Compton machine. Currently you can access this through the HACSBC2 workspace on adaql6.
If you can't find a VNC viewer open, or if it has crashed (usually with an error message saying the RECT is too large), you can open a new VNC viewer on any workspace with the command
The password is the same as the password for the adaq account. There is no username.
In the VNC window, this runcontrol window is in the workspace labeled "HAPPEX CODA" -- second column, bottom row. We are running in the "d2n" configuration.
Compton Counting DAQ
In the VNC window, this runcontrol window is in the workspace labeled CODA -- second column, top row. If Alexandre or Sirish is around, check with them before stopping and starting a run. The configuration is "Production."
Compton FADC DAQ
In the VNC window, this runcontrol window is in the workspace labeled CMU CODA -- next-to-last column, bottom row. If Diana is around, check with her before stopping and starting a run. The configuration is "CMUAccum."